# GPU-Accelerated Counterfactual Regret Minimization

### v4

@@ Abstract
Counterfactual regret minimization is a family of algorithms of no-regret learning dynamics capable of solving large-scale imperfect information games. We propose implementing this algorithm as a series of dense and sparse matrix and vector operations, thereby making it highly parallelizable for a graphical processing unit, at a cost of higher memory usage. Our experiments show that our implementation performs up to about 244.5 times faster than OpenSpiel’s Python implementation and, on an expanded set of games, up to about 114.2 times faster than OpenSpiel’s C++ implementation and the speedup becomes more pronounced as the size of the game being solved grows.
@@ 1 Introduction
Our experiments show that, compared to Google DeepMind OpenSpiel’s (Lanctot et al., 2020) Python implementation, our GPU implementation performs about 3.5 times slower for small games but is up to about 244.5 times faster for large games. Against their C++ implementation, our performance with a GPU is up to about 85.2 times slower for small games, but, on an expanded set of games, is up to about 114.2 times faster for large games. Even without a GPU, our implementation shows speedups compared to the OpenSpiel baselines (from about 1.5 to 46.8 times faster than their Python implementation and from 16.8 times slower to 4.5 times faster than theirs in C++). We see that the speedup becomes more pronounced as the size of the game solved grows.
@@ 3.1 Setup
${\bm{M}}^{(H_{+},Q_{+})}=\left(\bm{1}_{h_{+}=h_{+}^{\prime}}\right)_{(h_{+},(h_{+}^{\prime},a))\in{\mathbb{H}}_{+}\times{\mathbb{Q}}_{+}}$
${\bm{M}}^{(V,I_{+})}=\left(\begin{cases}\bm{1}_{f_{i}(f_{h}(f_{Pa}(v)))=i_{+}}&v\in{\mathbb{V}}_{+}\\ 0&v=v_{0}\\ \end{cases}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}$
@@ 3.2.1 Tree Traversal
$\widecheck{{\bm{U}}}^{(D+1)}=\left(\begin{cases}u(v,i_{+})&v\in{\mathbb{T}}\\ 0&v\in{\mathbb{D}}\end{cases}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}$
$\forall l\in[1,D]\cap{\mathbb{Z}}\quad\widecheck{{\bm{U}}}^{(l)}=\left({\bm{L}}^{(l)}\odot{\bm{S}}\right)\widecheck{{\bm{U}}}^{(l+1)}+\widecheck{{\bm{U}}}^{(l+1)}$
@@
$\widecheck{{\bm{\Pi}}}^{(0)}=\left(\bm{1}_{v=v_{0}}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}$
$\forall l\in[1,D]\cap{\mathbb{Z}}\quad\widecheck{{\bm{\Pi}}}^{(l)}=\left(\left({\bm{L}}^{(l)}\right)^{\top}\widecheck{{\bm{\Pi}}}^{(l-1)}\right)\odot\widecheck{{\bm{S}}}+\widecheck{{\bm{\Pi}}}^{(l-1)}$ (13)
@@ 4.1 Experiment 1
In our GPU implementation (written in Python), we use CuPy (Okuta et al., 2017) for GPU-accelerated matrix and vector operations. For parity with OpenSpiel (Lanctot et al., 2020), our implementation uses double-precision floating point numbers (64-bit float) and does not leverage tensor cores. We also simply run our implementation with NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020) (i.e., without a GPU) which we refer to as our CPU implementation. Our testbench computer contains an AMD Ryzen 9 3900X 12-core, 24-thread desktop processor, 128 GB memory, and Nvidia GeForce RTX 4090 24 GB VRAM graphics card.
@@ 4.1.2 Medium Games: Kuhn Poker (3-Player), First Sealed Auction, and Leduc Poker
In medium-sized games like Kuhn poker (3-player) (617 nodes), first sealed auction (7,096 nodes), and Leduc poker (9,457 nodes), performance gains compared to OpenSpiel’s (Lanctot et al., 2020) Python implementation can be observed for both our CPU (about 12.9, 46.8, and 44.6 times faster, respectively) and GPU implementation (about 3.2, 24.2, and 24.5 times faster, respectively). However, comparisons with OpenSpiel’s C++ implementation are mixed. For Kuhn poker (3-player), OpenSpiel’s C++ implementation is about 1.6 times faster than our CPU implementation and 6.5 times faster than our GPU implementation. But, for first sealed auction and Leduc poker, our CPU implementation is about 2.1 and 4.5 times faster, respectively, and our GPU implementation is about 1.1 and 2.5 times faster, respectively, than their C++ baseline. Here, while we begin to see our implementations outperform OpenSpiel’s baselines, we see that our CPU implementation is faster than our GPU implementation. This suggests that, while the efficiency of our implementation overcomes the Python overhead, the remaining GPU overhead makes using a GPU less preferable than not.
@@ 4.1.3 Large Games: Tiny Bridge (2-Player), Liar’s Dice, and Tic-Tac-Toe
In games like tiny bridge (2-player) (107,129 nodes), liar’s dice (294,883 nodes), and tic-tac-toe (549,946 nodes), noticeable performance gains over OpenSpiel’s (Lanctot et al., 2020) Python implementation can be observed for both our CPU (about 26.1, 13.9, and 16.8 times faster, respectively) and GPU implementation (about 108.6, 125.5, and 244.5 times faster, respectively). The same can be said for OpenSpiel’s C++ implementation to a lesser degree: our CPU implementation is about 1.5, 1.0, and 1.1 times faster, respectively, and our GPU implementation is about 6.4, 9.1, and 15.4 times faster, respectively. Here, the performance benefits of utilizing a GPU are clear, and we predict that the differences will be even more pronounced for games of sizes larger than the ones explored.
@@ 4.1.4 Memory Usages
Table 1: The peak memory usage of the benchmark scripts of the 4 CFR implementations. For our GPU implementation, we show both the peak memory usage of the benchmark process and the total memory allocated by CuPy in the CUDA memory pool.
@@ 4.2 Experiment 2
In order to see further scaling behavior or our GPU and OpenSpiel’s C++ implementations, we run these implementations with much larger imperfect information games (12 battleship games of differing parameters) for 10 iterations each. The iteration times for Experiment 2 are tabulated in Appendix B. For games as large as up to over 57.9 million nodes, our GPU implementation performs up to 114.2 times faster than OpenSpiel’s C++ baseline. Again, the relationship between the game sizes and the runtimes of each implementation is shown more clearly in the log-log graph in Figure 1.
@@ Appendix A Experiment 1
tiny_hanabi 0.851 (0.00) 0.035 (0.00) 0.581 (0.00) 2.958 (0.10)
kuhn_poker 1.011 (0.00) 0.042 (0.00) 0.684 (0.00) 3.319 (0.01)
kuhn_poker(players=3) 15.224 (0.01) 0.725 (0.00) 1.177 (0.00) 4.692 (0.01)
first_sealed_auction 81.226 (0.02) 3.696 (0.01) 1.736 (0.00) 3.355 (0.01)
leduc_poker 153.731 (0.19) 15.444 (0.02) 3.449 (0.00) 6.269 (0.01)
tiny_bridge_2p 640.783 (1.57) 37.524 (0.25) 24.513 (0.02) 5.902 (0.01)
liars_dice 1351.281 (8.39) 98.109 (0.79) 96.939 (0.07) 10.766 (0.02)
tic_tac_toe 2629.924 (11.04) 165.389 (0.78) 156.429 (0.15) 10.756 (0.02)
@@
tiny_hanabi 1.5 -3.5 -16.8 -85.2
kuhn_poker 1.5 -3.3 -16.1 -78.3
kuhn_poker(players=3) 12.9 3.2 -1.6 -6.5
first_sealed_auction 46.8 24.2 2.1 1.1
leduc_poker 44.6 24.5 4.5 2.5
tiny_bridge_2p 26.1 108.6 1.5 6.4
liars_dice 13.9 125.5 1.0 9.1
tic_tac_toe 16.8 244.5 1.1 15.4
@@
tiny_hanabi 0.031
kuhn_poker 0.032
kuhn_poker(players=3) 0.261
first_sealed_auction 1.911
leduc_poker 3.224
tiny_bridge_2p 35.327
liars_dice 104.731
tic_tac_toe 217.263
The total allocated CUDA memory by CuPy (Okuta et al., 2017) in our GPU implementation to solve each game through CFR in Experiment 1 is plotted in Figure 2 and tabulated in Table 4.
@@
While exploitability is a concept only valid for 2-player zero-sum games (and our method is also applicable to and tested on non-2-player general-sum games), we ran our algorithm again (separately from the benchmarks) with exploitabilities for Experiment 1. These are plotted in Figure 3. Note that the convergence behavior of CFR is already well-known, and our contribution is not about optimizing the convergence metrics like exploitability as most past CFR works have been about. We calculated exploitabilities purely for the sanity testing of our implementation.
@@ Appendix B Experiment 2
In Experiment 2, we tested two implementations – OpenSpiel’s C++ and our GPU – in 12 very large battleship games (up to over 57.9 million nodes) by running CFR for 10 iterations.
@@
Battleship-0 2,2,[1],2 5.134 (0.26) 14.197 (10.21)
Battleship-1 2,2,[1;2],2 61.807 (4.50) 5.099 (0.20)
Battleship-2 2,2,[1],3 40.392 (4.21) 5.247 (0.21)
Battleship-3 2,3,[1],2 59.662 (3.53) 4.104 (0.27)
Battleship-4 2,2,[1;2],3 959.262 (122.18) 9.558 (0.42)
Battleship-5 3,3,[1],2 676.297 (36.66) 17.034 (9.80)
Battleship-6 2,3,[1],3 1499.620 (120.59) 15.913 (0.83)
Battleship-7 3,4,[1],2 4161.539 (159.38) 38.554 (2.36)
Battleship-8 4,4,[1],2 23262.634 (1034.90) 233.995 (9.59)
Battleship-9 2,3,[1],4 33245.108 (4039.16) 528.117 (25.22)
Battleship-10 3,3,[1;2],2 106613.962 (6977.96) 933.899 (48.55)
Battleship-11 4,5,[1],2 90346.083 (5783.18) 856.541 (59.31)
@@
Battleship-0 -2.8
Battleship-1 12.1
Battleship-2 7.7
Battleship-3 14.5
Battleship-4 100.4
Battleship-5 39.7
Battleship-6 94.2
Battleship-7 107.9
Battleship-8 99.4
Battleship-9 63.0
Battleship-10 114.2
Battleship-11 105.5
The raw values and speedups (or slowdowns) are tabulated in Table 5 and Table 6, respectively. Note that our GPU implementation performs about 2.8 times slower for the zeroth battleship game, but is up to 114.2 times faster (for the tenth battleship game).

### v5

@@ Abstract
Counterfactual regret minimization is a family of algorithms of no-regret learning dynamics capable of solving large-scale imperfect information games. We propose implementing this algorithm as a series of dense and sparse matrix and vector operations, thereby making it highly parallelizable for a graphical processing unit, at a cost of higher memory usage. Our experiments show that our implementation performs up to about 401.2 times faster than OpenSpiel’s Python implementation and, on an expanded set of games, up to about 203.6 times faster than OpenSpiel’s C++ implementation and the speedup becomes more pronounced as the size of the game being solved grows.
@@ 1 Introduction
Our experiments show that, compared to Google DeepMind OpenSpiel’s (Lanctot et al., 2020) Python implementation, our GPU implementation performs about 3.5 times slower for small games but is up to about 401.2 times faster for large games. Against their C++ implementation, our performance with a GPU is up to about 85.5 times slower for small games, but, on an expanded set of games, is up to about 203.6 times faster for large games. Even without a GPU, our implementation shows speedups compared to the OpenSpiel baselines (from about 1.5 to 46.8 times faster than their Python implementation and from 16.8 times slower to 4.5 times faster than theirs in C++). We see that the speedup becomes more pronounced as the size of the game solved grows.
@@ 3.1 Setup
${\bm{M}}^{(H_{+},Q_{+})}=\left(\bm{1}_{h_{+}=h_{+}^{\prime}}\right)_{(h_{+},(h_{+}^{\prime},a))\in{\mathbb{H}}_{+}\times{\mathbb{Q}}_{+}}\qquad{\bm{M}}^{(V,I_{+})}=\left(\begin{cases}\bm{1}_{f_{i}(f_{h}(f_{Pa}(v)))=i_{+}}&v\in{\mathbb{V}}_{+}\\ 0&v=v_{0}\\ \end{cases}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}$
@@ 3.2.1 Tree Traversal
$\widecheck{{\bm{U}}}^{(D+1)}=\left(\begin{cases}u(v,i_{+})&v\in{\mathbb{T}}\\ 0&v\in{\mathbb{D}}\end{cases}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}\qquad\forall l\in[1,D]\cap{\mathbb{Z}}\quad\widecheck{{\bm{U}}}^{(l)}=\left({\bm{L}}^{(l)}\odot{\bm{S}}\right)\widecheck{{\bm{U}}}^{(l+1)}+\widecheck{{\bm{U}}}^{(l+1)}$
@@
$\widecheck{{\bm{\Pi}}}^{(0)}=\left(\bm{1}_{v=v_{0}}\right)_{(v,i_{+})\in{\mathbb{V}}\times{\mathbb{I}}_{+}}\qquad\forall l\in[1,D]\cap{\mathbb{Z}}\quad\widecheck{{\bm{\Pi}}}^{(l)}=\left(\left({\bm{L}}^{(l)}\right)^{\top}\widecheck{{\bm{\Pi}}}^{(l-1)}\right)\odot\widecheck{{\bm{S}}}+\widecheck{{\bm{\Pi}}}^{(l-1)}$
@@ 4.1 Experiment 1
In our GPU implementation (written in Python), we use CuPy (Okuta et al., 2017) for GPU-accelerated matrix and vector operations with both 64-bit and 32-bit floating-point data types. Here, we only discuss the 32-bit implementation as it is generally faster than the 64-bit version. Note that we save both memory and runtime if single-precision floating point numbers are used (by roughly a factor of 2). We also simply run our implementation with NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020) (i.e., without a GPU) which we refer to as our CPU implementation (with 64-bit floats). Our testbench computer contains an AMD Ryzen 9 3900X 12-core, 24-thread desktop processor, 128 GB memory, and Nvidia GeForce RTX 4090 24 GB VRAM graphics card.
@@ 4.1.2 Medium Games: Kuhn Poker (3-Player), First Sealed Auction, and Leduc Poker
In medium-sized games like Kuhn poker (3-player) (617 nodes), first sealed auction (7,096 nodes), and Leduc poker (9,457 nodes), performance gains compared to OpenSpiel’s (Lanctot et al., 2020) Python implementation can be observed for both our CPU (about 12.9, 46.8, and 44.6 times faster, respectively) and GPU implementation (about 3.1, 23.2, and 24.9 times faster, respectively). However, comparisons with OpenSpiel’s C++ implementation are mixed. For Kuhn poker (3-player), OpenSpiel’s C++ implementation is about 1.6 times faster than our CPU implementation and 6.8 times faster than our GPU implementation. But, for first sealed auction and Leduc poker, our CPU implementation is about 2.1 and 4.5 times faster, respectively, and our GPU implementation is about 1.1 and 2.5 times faster, respectively, than their C++ baseline. Here, while we begin to see our implementations outperform OpenSpiel’s baselines, we see that our CPU implementation is faster than our GPU implementation. This suggests that, while the efficiency of our implementation overcomes the Python overhead, the remaining GPU overhead makes using a GPU less preferable than not.
@@ 4.1.3 Large Games: Tiny Bridge (2-Player), Liar’s Dice, and Tic-Tac-Toe
In games like tiny bridge (2-player) (107,129 nodes), liar’s dice (294,883 nodes), and tic-tac-toe (549,946 nodes), noticeable performance gains over OpenSpiel’s (Lanctot et al., 2020) Python implementation can be observed for both our CPU (about 26.1, 13.9, and 16.8 times faster, respectively) and GPU implementation (about 111.8, 160.0, and 401.2 times faster, respectively). The same can be said for OpenSpiel’s C++ implementation to a lesser degree: our CPU implementation is about 1.5, 1.0, and 1.1 times faster, respectively, and our GPU implementation is about 6.5, 11.6, and 25.2 times faster, respectively. Here, the performance benefits of utilizing a GPU are clear, and we predict that the differences will be even more pronounced for games of sizes larger than the ones explored.
@@ 4.1.4 Memory Usages
Table 1: The peak memory usage of the benchmark scripts of the 4 CFR implementations. For GPU implementations, peak usages of the process and the memory allocated by CuPy are shown.
@@
32-bit Process 2.371
CUDA 0.176
@@ 4.2 Experiment 2
In order to see further scaling behavior or our GPU (both 64-bit and 32-bit versions) and OpenSpiel’s C++ implementations, we run these implementations with much larger imperfect information games (12 battleship games of differing parameters) for 10 iterations each. The iteration times for Experiment 2 are tabulated in Appendix B. For games as large as up to over 57.9 million nodes, our GPU implementation performs up to 203.6 times faster than OpenSpiel’s C++ baseline. Again, the relationship between the game sizes and the runtimes is shown in the log-log graph in Figure 1.
@@ Appendix A Experiment 1
64-bit 32-bit
tiny_hanabi 0.851 (0.00) 0.035 (0.00) 0.581 (0.00) 2.958 (0.10) 2.968 (0.10)
kuhn_poker 1.011 (0.00) 0.042 (0.00) 0.684 (0.00) 3.319 (0.01) 3.362 (0.00)
kuhn_poker(players=3) 15.224 (0.01) 0.725 (0.00) 1.177 (0.00) 4.692 (0.01) 4.906 (0.01)
first_sealed_auction 81.226 (0.02) 3.696 (0.01) 1.736 (0.00) 3.355 (0.01) 3.495 (0.00)
leduc_poker 153.731 (0.19) 15.444 (0.02) 3.449 (0.00) 6.269 (0.01) 6.178 (0.01)
tiny_bridge_2p 640.783 (1.57) 37.524 (0.25) 24.513 (0.02) 5.902 (0.01) 5.732 (0.01)
liars_dice 1351.281 (8.39) 98.109 (0.79) 96.939 (0.07) 10.766 (0.02) 8.443 (0.01)
tic_tac_toe 2629.924 (11.04) 165.389 (0.78) 156.429 (0.15) 10.756 (0.02) 6.556 (0.00)
@@
64-bit 32-bit 64-bit 32-bit
tiny_hanabi 1.5 -3.5 -3.5 -16.8 -85.2 -85.5
kuhn_poker 1.5 -3.3 -3.3 -16.1 -78.3 -79.3
kuhn_poker(players=3) 12.9 3.2 3.1 -1.6 -6.5 -6.8
first_sealed_auction 46.8 24.2 23.2 2.1 1.1 1.1
leduc_poker 44.6 24.5 24.9 4.5 2.5 2.5
tiny_bridge_2p 26.1 108.6 111.8 1.5 6.4 6.5
liars_dice 13.9 125.5 160.0 1.0 9.1 11.6
tic_tac_toe 16.8 244.5 401.2 1.1 15.4 25.2
@@
Double-Precision (64-bit) Single-Precision (32-bit)
tiny_hanabi 0.031 0.026
kuhn_poker 0.032 0.028
kuhn_poker(players=3) 0.261 0.157
first_sealed_auction 1.911 1.121
leduc_poker 3.224 1.978
tiny_bridge_2p 35.327 20.373
liars_dice 104.731 65.520
tic_tac_toe 217.263 139.977
The total allocated CUDA memory by CuPy (Okuta et al., 2017) in our GPU implementation to solve each game through CFR in Experiment 1 is plotted in Figure 2 and tabulated in Table 4. Note that noticeable improvement in memory usage is seen when 32-bit floating-point numbers are used instead of 64-bit floating-point numbers.
@@
While exploitability is a concept only valid for 2-player zero-sum games (and our method is also applicable to and tested on non-2-player general-sum games), we ran our algorithm again (separately from the benchmarks) with exploitabilities for Experiment 1 (on GPU 64-bit implementation). These are plotted in Figure 3. Note that the convergence behavior of CFR is already well-known, and our contribution is not about optimizing the convergence metrics like exploitability as most past CFR works have been about. We calculated exploitabilities purely for the sanity testing of our implementation.
@@ Appendix B Experiment 2
In Experiment 2, we tested two implementations – OpenSpiel’s C++ and our GPU – in 12 very large battleship games (up to over 57.9 million nodes) by running CFR for 10 iterations. As these games are large, we loaded the game without storing all state objects into memory (instead, we use Python integers).
@@
64-bit 32-bit
Battleship-0 2,2,[1],2 5.134 (0.26) 14.197 (10.21) 14.393 (10.57)
Battleship-1 2,2,[1;2],2 61.807 (4.50) 5.099 (0.20) 4.880 (0.14)
Battleship-2 2,2,[1],3 40.392 (4.21) 5.247 (0.21) 4.939 (0.13)
Battleship-3 2,3,[1],2 59.662 (3.53) 4.104 (0.27) 3.983 (0.13)
Battleship-4 2,2,[1;2],3 959.262 (122.18) 9.558 (0.42) 7.094 (0.32)
Battleship-5 3,3,[1],2 676.297 (36.66) 17.034 (9.80) 6.510 (0.20)
Battleship-6 2,3,[1],3 1499.620 (120.59) 15.913 (0.83) 10.156 (0.43)
Battleship-7 3,4,[1],2 4161.539 (159.38) 38.554 (2.36) 20.437 (1.04)
Battleship-8 4,4,[1],2 23262.634 (1034.90) 233.995 (9.59) 131.795 (3.48)
Battleship-9 2,3,[1],4 33245.108 (4039.16) 528.117 (25.22) 384.525 (5.81)
Battleship-10 3,3,[1;2],2 106613.962 (6977.96) 933.899 (48.55) 563.536 (9.90)
Battleship-11 4,5,[1],2 90346.083 (5783.18) 856.541 (59.31) 446.369 (10.94)
@@
Double-Precision (64-bit) Single-Precision (32-bit)
Battleship-0 -2.8 -2.8
Battleship-1 12.1 12.7
Battleship-2 7.7 8.2
Battleship-3 14.5 15.0
Battleship-4 100.4 135.2
Battleship-5 39.7 103.9
Battleship-6 94.2 147.7
Battleship-7 107.9 203.6
Battleship-8 99.4 176.5
Battleship-9 63.0 86.5
Battleship-10 114.2 189.2
Battleship-11 105.5 202.4
The raw values and speedups (or slowdowns) are tabulated in Table 5 and Table 6, respectively. Note that our 32-bit GPU implementation performs about 2.8 times slower for the zeroth battleship game, but is up to 203.6 times faster (for the seventh battleship game).